OpenStudy (anonymous):

After solving this equation: 2x²+14x-60=0 What is the negative value of x?

OpenStudy (anonymous):

2x^2+14x-60=0 factor out the left side to so the equation equals: (2x-6)(x+10)=0 since (2x-6) will produce a positive x, find the value for (x+10): x+10=0 x=-10